### v2

# MagicTailor: Component-Controllable Personalization in Text-to-Image Diffusion Models

Abstract
* Equal contribution. $\dagger$ Corresponding author.
Recent text-to-image models generate high-quality images from text prompts but lack precise control over specific components within visual concepts. Therefore, we introduce component-controllable personalization, a new task that allows users to customize and reconfigure individual components within concepts. This task faces two challenges: semantic pollution, where undesirable elements distort the concept, and semantic imbalance, which leads to disproportionate learning of the target concept and component. To address these, we design MagicTailor, a framework that uses Dynamic Masked Degradation to adaptively perturb unwanted visual semantics and Dual-Stream Balancing for more balanced learning of desired visual semantics. The experimental results show that MagicTailor outperforms existing methods in this task and enables more personalized, nuanced, and creative image generation.
Figure 1: (a) Personalization Illustration: Showing how text-to-image diffusion models learn and generate a visual concept from reference images. (b) Component-Controllable Personalization Illustration: Introducing a task that aims to modify specific components of visual concepts during personalization. (c) Generated Images by MagicTailor: Demonstrating MagicTailor’s effectiveness in component-controllable personalization, with red and blue circles highlighting the target concept and component, respectively.
1 Introduction
Text-to-image (T2I) diffusion models [35, 39, 31, 5] have shown impressive capabilities in generating high-quality images from textual descriptions. While these models can generate images that align well with provided prompts, they struggle when certain visual concepts are hard to express in natural language. To address this, methods like [11, 36] enable T2I models to learn specific concepts from a few reference images, allowing for more accurate integration of those concepts into the generated images. This process, as shown in Fig. 1 (a), is referred as personalization.
However, existing personalization methods are limited to replicating predefined concepts and lack flexible and fine-grained control over their components. This limitation hinders their practical use in real-world applications. Since concepts often consist of multiple components, such as a house with walls, windows, and doors, a key problem in personalization is how to effectively control and manipulate these individual components.
Figure 2: Major challenges in component-controllable personalization. (a) Semantic pollution: (i) Undesired elements may interfere with the personalized concept. (ii) A simple mask-out strategy causes unintended results, while (iii) DM-Deg effectively suppresses unwanted semantics. (b) Semantic imbalance: (i) Simultaneously learning the concept and component can distort either one. (ii) DS-Bal ensures balanced learning, improving personalization. (c) Identity fidelity performance: Calculating DreamSim [10] scores on a collected dataset, we show that DM-Deg and DS-Bal can address these challenges for faithful generation.
In this paper, we introduce component-controllable personalization, a new task that enables the reconfiguration of specific components within personalized concepts using additional visual references (Fig. 1 (b)). In this approach, a T2I model is fine-tuned with reference images and corresponding category labels, allowing it to learn and generate the desired concept along with the given component. This capability empowers users to refine and customize concepts with precise control, fostering creativity and innovation across various domains, from artworks to inventions.
One key challenge of this task is semantic pollution (Fig. 2 (a)), where unwanted visual elements inadvertently appear in the generated image, “polluting” the personalized concept. This happens because the T2I model often mixes visual semantics from different regions during training. Masking out unwanted elements in reference images doesn’t solve the problem, as it disrupts the visual context and creates unintended compositions. Another challenge is semantic imbalance (Fig. 2 (b)), where the model overemphasizes certain aspects, leading to unfaithful personalization. This occurs due to the semantic disparity between the concept and component, necessitating a more balanced learning approach to manage concept-level (e.g., person) and component-level (e.g., hair) semantics.
Figure 3: Pipeline overview of MagicTailor. This method fine-tunes a T2I diffusion model using reference images to learn both the target concept and component, enabling the generation of images that seamlessly integrate the component into the concept. Two key techniques, Dynamic Masked Degradation (DM-Deg, see Sec. 3.2) and Dual-Stream Balancing (DS-Bal, see Sec. 3.3), address semantic pollution and semantic imbalance, respectively. For clarity, only one image per concept/component is shown, and the warm-up stage is omitted.
To address these challenges, we present MagicTailor, a novel framework that enables component-controllable personalization for T2I models ( Fig. 1 (c)). We first use a text-guided image segmenter to generate segmentation masks for both the concept and component and then design Dynamic Masked Degradation (DM-Deg) to transform reference images into randomly degraded versions, perturbing undesired visual semantics. This method helps suppress the model’s sensitivity to irrelevant details while preserving the overall visual context, effectively mitigating semantic pollution. Next, we initiate a warm-up phase for the T2I model, training it on the degraded images using a masked diffusion loss to focus on the desired semantics and a cross-attention loss to strengthen the correlation between these semantics and pseudo-words. To address semantic imbalance, we develop Dual-Stream Balancing (DS-Bal), a dual-stream learning paradigm that balances the learning of visual semantics. In this phase, the online denoising U-Net performs sample-wise min-max optimization, while the momentum denoising U-Net applies selective preservation regularization. This ensures more accurate and faithful personalization of the target concept and component.
We validate the effectiveness of MagicTailor through various qualitative and quantitative experiments, demonstrating state-of-the-art (SOTA) performance in component-controllable personalization. Detailed ablation studies further confirm the impact of the key techniques integrated into MagicTailor. Additionally, we show its potential for enabling a wide range of creative applications.
2 Related Works
Text-to-Image Generation.
Text-to-image (T2I) generation has made remarkable advancements in recent years, enabling the synthesis of vivid and diverse imagery based on textual descriptions. Early methods employed Generative Adversarial Networks (GANs) [33, 53, 28, 59], and transformers [6, 30, 7, 55, 4] also showed the potential in conditional generation. The advent of diffusion models has ushered in a new era in T2I generation [20, 39, 31, 5, 54]. Leveraging these models, a range of related applications has rapidly emerged, including image editing [20, 24, 16, 9, 3], image completion and translation [50, 49, 22], and controllable image generation [56, 45, 58]. Despite advancements in T2I diffusion models, generating images that accurately reflect specific, user-defined concepts remains a challenge. This study explores component-controllable personalization, which allows precise adjustment of specific concepts’ components using visual references.
Personalization.
Personalization seeks to adapt T2I models to generate given concepts using reference images. Initial approaches [11, 36] addressed this task by either optimizing text embeddings or fine-tuning the entire T2I model. Additionally, low-rank adaptation (LoRA) [15] has been widely adopted by the research community for personalization [37], offering an efficient and lightweight solution. The scope of personalization has expanded to encompass multiple concepts [19, 2, 13, 12, 25]. Besides, several studies [48, 21, 40, 46, 44] have explored tuning-free approaches for personalization, but these often necessitate training an encoder on large, domain-specific image datasets [57]. In this work, MagicTailor adopts the widely used paradigm of test-time optimization, providing stable performance and precise control over personalization.
3 Methodology
Let ={($\mathcal{I}=\{(\{I_{nk}\}_{k=1}^{K},c_{n})\}_{n=1}^{N}$ represents a concept-component pair with $N$ samples of concepts and components, where each sample contains $K$ reference images $\{I_{nk}\}_{k=1}^{K}$ with a category label $c_{n}$. In this work, we focus on a practical setting involving one concept and one component. Specifically, we set $N=2$ and define the first sample as a concept (e.g., dog) while the second one as a component (e.g., ear). In addition, these samples are associated with the pseudo-words $\mathcal{P}=\{p_{n}\}_{n=1}^{N}$ serving as their text identifiers. The goal of component-controllable personalization is to fine-tune a text-to-image (T2I) model to accurately learn both the concept and the component from $\mathcal{I}$. Using text prompts with $\mathcal{P}$, the fine-tuned model should generate images that integrate the personalized concept with the specified component.
This section begins by providing an overview of the MagicTailor pipeline in Sec. 3.1 and then delves into its two core techniques: Dynamic Masked Degradation in Sec. 3.2 and Dual-Stream Balancing in Sec. 3.3.
3.1 Overall Pipeline
The overall pipeline of MagicTailor is illustrated in Fig. 3. The process begins with identifying the desired concept or component within each reference image $I_{nk}$, employing an off-the-shelf text-guided image segmenter to generate a segmentation mask $M_{nk}$ based on $I_{nk}$ and its associated category label $c_{n}$. Conditioned on $M_{nk}$, we design Dynamic Masked Degradation (DM-Deg) to perturb undesired visual semantics within $I_{nk}$, addressing semantic pollution. At each training step, DM-Deg transforms $I_{nk}$ into a randomly degraded image $\hat{I}_{nk}$, with the degradation intensity being dynamically regulated. Subsequently, these degraded images, along with structured text prompts, are used to fine-tune a T2I diffusion model to facilitate concept and component learning. The model is formally expressed as $\{\epsilon_{\theta},\tau_{\theta},\mathcal{E},\mathcal{D}\}$, where $\epsilon_{\theta}$ represents the denoising U-Net, $\tau_{\theta}$ is the text encoder, and $\mathcal{E}$ and $\mathcal{D}$ denote the image encoder and decoder, respectively. To promote the learning of the desired visual semantics, we employ the masked diffusion loss, which is defined as:
$\mathcal{L}_{\text{diff}}\ =\ \mathbb{E}_{n,k,\epsilon,t}\Big{[}\big{\|}\epsilon_{n}\odot M_{nk}^{\prime}-\epsilon_{\theta}(z_{nk}^{(t)},t,e_{n})\odot M_{nk}^{\prime}\big{\|}_{2}^{2}\Big{]}\ ,$ (1)
where $\epsilon_{n}\sim\mathcal{N}(0,1)$ is the unscaled noise, $z_{nk}^{(t)}$ is the noisy latent image of $\hat{I}_{nk}$ with a random time step $t$, $e_{n}$ is the text embedding of the corresponding text prompt, and $M_{nk}^{\prime}$ is downsampled from $M_{nk}$ to match the shape of $\epsilon$ and $z_{nk}$. Additionally, we incorporate the cross-attention loss to strengthen the correlation between desired visual semantics and their corresponding pseudo-words, formulated as:
$\mathcal{L}_{\text{attn}}\ =\ \mathbb{E}_{n,k,t}\Big{[}\big{\|}A_{\theta}(p_{n},z_{nk}^{(t)})-M_{nk}^{\prime\prime}\big{\|}_{2}^{2}\Big{]}\ ,$ (2)
when $A_{\theta}(p_{n},z_{nk}^{(t)})$ is the cross-attention maps between the pseudo-word $p_{n}$ and the noisy latent image $z_{nk}^{(t)}$ and $M_{nk}^{\prime\prime}$ is downsampled from $M_{nk}$ to match the shape of $A_{\theta}(p_{n},z_{nk}^{(t)})$.
Figure 4: Motivation of dynamic intensity. (a) Fixed intensity ($\alpha_{d}=0.5$ here) could cause noisy generated images. (b) Our dynamic intensity can mitigate noise memorization. (c) We report IQA results of Q-Align [47] on our dataset, showing that our dynamic intensity helps to enhance the quality of generated images.
Using $\mathcal{L}_{\text{diff}}$ and $\mathcal{L}_{\text{attn}}$, we first warm up the T2I model by jointly learning all samples, aiming to preliminarily inject the knowledge of visual semantics. The loss of the warm-up stage is defined as:
$\mathcal{L}_{\text{warm-up}}\ =\ \mathcal{L}_{\text{diff}}\ +\ \lambda_{\text{attn}}\mathcal{L}_{\text{attn}}\ ,$ (3)
where $\lambda_{\text{attn}}=0.01$ is the loss weight for $\mathcal{L}_{\text{attn}}$. For efficient fine-tuning, we only train the denoising U-Net $\epsilon_{\theta}$ in a low-rank adaptation (LoRA) [15] manner and the text embedding of the pseudo-words $\mathcal{P}$, keeping the others frozen. Thereafter, we employ Dual-Stream Balancing (DS-Bal) to address semantic imbalance. In this paradigm, the online denoising U-Net $\epsilon_{\theta}$ conducts sample-wise min-max optimization for the hardest-to-learn sample, and meanwhile the momentum denoising U-Net $\tilde{\epsilon}_{\theta}$ applies selective preserving regularization for the other samples.
3.2 Dynamic Masked Degradation
Semantic pollution is a major challenge for component-controllable personalization. As shown in Fig. 2 (a.i), the target concept (i.e., person) can be distorted by the owner of the target component (i.e., eye), resulting in a hybrid person. Masking regions outside the target concept and component can damage the overall context, leading to overfitting and odd compositions (Fig. 2 (a.ii)). To address this, undesired visual semantics in reference images must be handled appropriately. We propose Dynamic Masked Degradation (DM-Deg), which dynamically perturbs undesired semantics to suppress their influence on the T2I model while preserving the overall visual context (Fig. 2 (a.iii)&(c)).
Degradation Imposition.
In each training step, DM-Deg imposes degradation in the out-of-mask region for each reference image. We use Gaussian noise for degradation due to its simplicity. For a reference image $I_{nk}$, we randomly sample a Gaussian noise matrix $G_{nk}\sim\mathcal{N}(0,1)$ with the same shape as $I_{nk}$, where the pixel values of $I_{nk}$ range from $-1$ to $1$. The degradation is then applied as follows:
$\hat{I}_{nk}=\alpha_{d}G_{nk}\odot(1-M_{nk})+I_{nk},$ (4)
where $\odot$ denotes element-wise multiplication, and $\alpha_{d}\in[0,1]$ is a dynamic weight controlling the degradation intensity. While previous works [48, 21] have used noise to fully cover the background or enhance data diversity, DM-Deg aims to produce a degraded image $\hat{I}_{nk}$ that retains the original visual context. By introducing $\hat{I}_{nk}$, we can suppress the T2I model from perceiving undesired visual semantics in out-of-mask regions, as these semantics are perturbed by random noise at each training step.
Dynamic Intensity.
Unfortunately, the T2I model may gradually memorize the introduced noise while learning meaningful visual semantics, leading to noise appearing in generated images (Fig. 4 (a)). This behavior is consistent with previous observations on deep networks [1]. To address this, we propose a descending scheme that dynamically regulates the intensity of the imposed noise during training. This scheme follows an exponential curve, maintaining a relatively high intensity in the early stages and decreasing it sharply in later stages. Let $d$ denote the current training step and $D$ denote the total training step. The curve of dynamic intensity is defined as:
$\alpha_{d}\ =\ \alpha_{\text{init}}(1-(\frac{d}{D})^{\gamma})\ ,$ (5)
where $\alpha_{\text{init}}$ is the initial value of $\alpha_{d}$ and $\gamma$ controls the descent rate. We empirically set $\alpha_{\text{init}}=0.5$ and $\gamma=32$, tuned within the powers of $2$. This dynamic intensity scheme effectively prevents semantic pollution and mitigates the memorization of introduced noise, leading to improved generation performance (Fig. 4 (b)&(c)).
3.3 Dual-Stream Balancing
Another challenge is semantic imbalance, which arises from the disparity in visual semantics between the target concept and its component. Concepts generally have richer visual semantics than components (e.g. person vs. hair), but in some cases, the component may have more complex semantics (e.g., simple tower vs. intricate roof). This imbalance complicates joint learning, leading to overemphasis on either the concept or the component, and resulting in incoherent generation (Fig. 5 (a)). To address this, we design Dual-Stream Balancing (DS-Bal), a dual-stream learning paradigm with online and momentum denoising U-Nets (Fig. 3) for balanced semantic learning, improving personalization fidelity (Fig. 5 (b) & Fig. 2 (c)).
Figure 5: Learning process visualization. (a) The vanilla learning paradigm tends to overemphasize the easier one. (b) DS-Bal effectively balances the learning of the concept and component.
Sample-Wise Min-Max Optimization.
From a loss perspective, the visual semantics of the concept and component are learned by optimizing the masked diffusion loss $\mathcal{L}_{\text{diff}}$ across all the samples. However, this indiscriminate optimization fails to allocate sufficient learning effort to a more challenging sample, leading to an imbalanced learning process. To address this, DS-Bal uses the online denoising U-Net to focus on learning the hardest-to-learn sample at each training step. Inheriting the weights of the original denoising U-Net, which is warmed up through joint learning, the online denoising U-Net $\epsilon_{\theta}$ optimizes only the sample with the highest masked diffusion loss as:
$\mathcal{L}_{\text{diff-max}}\ =\ \max_{n}\mathbb{E}_{k,\epsilon,t}\Big{[}\big{\|}$ $\epsilon_{n}\odot M_{nk}^{\prime}-$
$\epsilon_{\theta}(z_{nk}^{(t)},t,e_{n})\odot M_{nk}^{\prime}\big{\|}_{2}^{2}\Big{]}\ ,$ (6)
where minimizing $\mathcal{L}_{\text{diff-max}}$ can be considered as a form of min-max optimization [32]. The learning objective of $\epsilon_{\theta}$ may switch across different training steps and is not consistently dominated by the concept or component. Such an optimization scheme can effectively modulate the learning dynamics of different samples and avoid the overemphasis on any particular one.
Selective Preserving Regularization.
At a training step, the sample neglected in $\mathcal{L}_{\text{diff-max}}$ may suffer from knowledge forgetting. This is because the optimization of $\mathcal{L}_{\text{diff-max}}$, which aims to enhance the knowledge of a specific sample, could inadvertently overshadow the knowledge of the others. In light of this, DS-Bal meanwhile exploits the momentum denoising U-Net to preserve the learned visual semantics of the other sample in each training step. Specifically, we first select the sample that is excluded in $\mathcal{L}_{\text{diff-max}}$, which is expressed as $S=\{n|n=1,...,N\}-\{n_{\text{max}}\},$ where $n_{\text{max}}$ is the index of the target sample in $\mathcal{L}_{\text{diff-max}}$ and $S$ is the selected index set. Then, we use the momentum denoising U-Net $\tilde{\epsilon}_{\theta}$ to apply regularization for $S$, with the masked preserving loss as:
$\mathcal{L}_{\text{pres}}\ =\ \mathbb{E}_{n\in S,k,t}\Big{[}\big{\|}\tilde{\epsilon}_{\theta}(z_{nk}^{(t)},t,e_{n})$ $\odot M_{nk}^{\prime}-$
$\epsilon_{\theta}(z_{nk}^{(t)},t,e_{n})$ $\odot M_{nk}^{\prime}\big{\|}_{2}^{2}\Big{]}\ ,$ (7)
where $\tilde{\epsilon}_{\theta}$ is updated from $\epsilon_{\theta}$ using EMA [42] with the smoothing coefficient $\beta=0.99$, thereby sustaining the prior accumulated knowledge of $\epsilon_{\theta}$ in each training step. By encouraging the consistency between the output of $\epsilon_{\theta}$ and $\tilde{\epsilon}_{\theta}$ in $\mathcal{L}_{\text{pres}}$, we can facilitate the knowledge maintenance of the other samples while learning a specific sample in $\mathcal{L}_{\text{diff-max}}$. Overall, DS-Bal can be considered a mechanism to adaptively assign target labels $\epsilon_{n}$ or preserving labels $\tilde{\epsilon}_{\theta}(z_{nk}^{(t)},t,e_{n})$ to different samples, enabling dynamic loss supervision (Fig. 3). Using a loss weight $\lambda_{\text{pres}}=0.2$, the total loss of the DS-Bal stage is formulated as:
$\mathcal{L}_{\text{DS-Bal}}\ =\ \mathcal{L}_{\text{diff-max}}\ +\ \lambda_{\text{pres}}\mathcal{L}_{\text{pres}}\ +\ \lambda_{\text{attn}}\mathcal{L}_{\text{attn}}\ .$ (8)
4 Experimental Results
Figure 6: Qualitative comparisons. We present images generated by MagicTailor and others across various domains. MagicTailor achieves strong text alignment, identity fidelity, and high-quality generation. More results are in the supplementary material.
4.1 Experimental Setup
Dataset, Implementation, and Evaluation.
For a systematic investigation, we compile a dataset from diverse domains, including characters, animation, buildings, objects, and animals. We use Stable Diffusion 2.1 [35] as the pretrained T2I model and resize all reference images to $512\times 512$. For the warm-up and DS-Bal stages, we set the training steps to 200 and 300, with learning rates of $1\times 10^{-4}$ and $1\times 10^{-5}$, respectively. Leveraging LoRA fine-tuning, each concept-component pair requires only about five minutes of training on an A100 GPU. For evaluation, we design 20 text prompts covering a wide range of scenarios and generate 14,720 images for each method. To ensure fairness, all random seeds are fixed during both training and inference.
Compared Methods.
We compare our MagicTailor with state-of-the-art (SOTA) personalization methods, including Textual Inversion (TI) [11], DreamBooth (DB) [36], Custom Diffusion (CD) [19], Break-A-Scene (BAS) [2], and CLiC [38]. These methods were selected for their representativeness of personalization frameworks or relevance to learning fine-grained elements. For a fair comparison, we adapt them to our task with minimal modifications, specifically by incorporating the masked diffusion loss (Eq. 1). Apart from method-specific configurations, all methods are implemented using the same setup to ensure consistency. Due to space constraints, additional details on the experimental setup can be found in the supplementary material.
4.2 Qualitative Comparisons
The qualitative results are shown in in Fig. 6. As observed, TI, CD, and CLiC primarily suffer from semantic pollution, where undesired visual semantics significantly distort the personalized concept. Besides, DB and BAS also struggle in this challenging task, with an overemphasis on either the concept or the component due to semantic imbalance, sometimes even causing the target component to be completely absent. An interesting finding is that imbalanced learning can exacerbate semantic pollution, leading to the color and texture of the target concept or component being mistakenly transferred to unintended parts of the generated images. In contrast, MagicTailor effectively generates text-aligned images that accurately represent both the target concept and component. To further validate its performance, we provide additional comparisons and a generalizability analysis in the supplementary material.
Table 1: Quantitative comparisons on automatic metrics. MagicTailor can achieve SOTA performance on all four automatic metrics. The best results are marked in bold.
Methods CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO $\uparrow$ DreamSim $\downarrow$
Textual Inversion [11] 0.236 0.742 0.620 0.558
DreamBooth [36] 0.266 0.841 0.798 0.323
Custom Diffusion [19] 0.251 0.797 0.750 0.407
Break-A-Scene [2] 0.259 0.840 0.780 0.338
CLiC [38] 0.263 0.764 0.663 0.499
MagicTailor (Ours) 0.270 0.854 0.813 0.279
Table 2: Quantitative comparisons on the user study. MagicTailor can also outperform other methods across all three aspects in human evaluation.
Methods Text Align. $\uparrow$ Id. Fidelity $\uparrow$ Gen. Quality $\uparrow$
Textual Inversion [11] 5.8% 2.5% 5.2%
DreamBooth [36] 15.3% 14.7% 12.5%
Custom Diffusion [19] 7.1% 7.7% 9.8%
Break-A-Scene [2] 10.8% 12.1% 22.8%
CLiC [38] 4.5% 5.1% 6.2%
MagicTailor (Ours) 56.5% 57.9% 43.4%
Table 3: Ablation of key techniques. Our DM-Deg and DS-Bal effectively contribute to a superior performance trade-off.
DM-Deg DS-Bal CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO $\uparrow$ DreamSim $\downarrow$
0.275 0.837 0.798 0.317
$\checkmark$ 0.276 0.848 0.809 0.294
$\checkmark$ 0.270 0.845 0.802 0.304
$\checkmark$ $\checkmark$ 0.270 0.854 0.813 0.279
4.3 Quantitative Comparisons
Automatic Metrics. We utilize four automatic metrics in the aspects of text alignment (CLIP-T [11]) and identity fidelity (CLIP-I [29], DINO [26], DreamSim [10]). To precisely meafsure identity fidelity, we segment out the concept and component in each reference and evaluation image, and then eliminate the target component from the segmented concept. As we can see in Tab. 1, component-controllable personalization remains a tough task even for SOTA methods of personalization. By comparison, MagicTailor can achieve the best results in both identity fidelity and text alignment. It should be credited to the utilization of an effective framework tailored to this special task.
User Study.
We further evaluate the methods with a user study. Specifically, a detailed questionnaire is designed to display 20 groups of evaluation images with the corresponding text prompt and reference images. Users are asked to select the best result in each group for three aspects, including text alignment, identity fidelity, and generation quality. Finally, we collect a total of 3,180 valid answers and report the selected rates in Tab. 2. It can be observed that MagicTailor can also achieve superior performance in human preferences, showcasing its effectiveness in component-controllable personalization.
Table 4: Ablation of DM-Deg. We compare DM-Deg with its variants and the mask-out strategy. Our DM-Deg attains superior overall performance on text alignment and identity fidelity.
Intensity Variants CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO $\uparrow$ DreamSim $\downarrow$
Mask-Out Startegy 0.270 0.818 0.760 0.375
Fixed ($\alpha=0.4$) 0.270 0.849 0.800 0.297
Fixed ($\alpha=0.6$) 0.271 0.845 0.794 0.310
Fixed ($\alpha=0.8$) 0.271 0.846 0.796 0.305
Linear (Ascent) 0.270 0.846 0.797 0.307
Linear (Descent) 0.261 0.851 0.802 0.300
Dynamic ($\gamma=8$) 0.266 0.850 0.806 0.289
Dynamic ($\gamma=16$) 0.268 0.854 0.813 0.282
Dynamic ($\gamma=64$) 0.271 0.852 0.812 0.283
Dynamic (Ours) 0.270 0.854 0.813 0.279
4.4 Ablation Studies
We conduct comprehensive ablation studies for MagicTailor to verify the capability of the overall pipeline. More ablation studies are included in the supplementary material.
Effectiveness of Key Techniques.
In Tab. 3, we investigate two key techniques of MagicTailor by starting from a baseline framework described in Sec. 3.1. Even without DM-Deg and DS-Bal, such a baseline framework can still have competitive performance, showing its reliability. On top of that, we introduce DM-Deg and DS-Bal, where the superior performance trade-off indicates their significance. Qualitative results can also refer to Fig. 2.
Figure 7: Applications of MagicTailor. (a) Decoupled generation. MagicTailor can also separately generate the target concept and component, enriching prospective combinations. (b) Controlling multiple components. MagicTailor shows the potential to handle more than one component, highlighting its effectiveness. (c) Enhancing other generative tools. MagicTailor can seamlessly integrate with various generative tools, adding the capability to control components within their generation pipelines.
Dynamic Intensity Matters.
In Tab. 4, we explore DM-Deg by comparing it with 1) mask-out strategy; 2) fixed intensity; 3) linear intensity ($\alpha$ goes from 1 to 0, or from 0 to 1); and 4) dynamic intensity with different $\gamma$. First, the terrible performance of the mask-out strategy verifies that it is not a valid solution for semantic pollution. Notably, the descent linear intensity shows better identity fidelity than its ascent counterpart, which aligns with and validates our observations on noise memorization. Moreover, the dynamic intensity generally shows better results, and it can achieve better overall performance with a proper $\gamma$.
Momentum Denoising U-Net as a Good Regularizer.
In Tab. 5, we study DS-Bal by modifying the U-Net for regularization as 1) fixed U-Net with $\beta=0$ (i.e., the one just after warm-up); 2) fixed U-Net with $\beta=1$ (i.e., the one from the last step); and 3) momentum U-Net with other $\beta$. The results show that employing the U-Net with a high momentum rate can yield better regularization to tackle semantic imbalance, thus leading to excellent performance.
Sensitivity Analysis of Loss Weights.
In Fig. 8, we analyze the sensitivity of loss weights in Eq. 8 (i.e., $\lambda_{\text{pres}}$ and $\lambda_{\text{attn}}$), since loss weights are often critical for model training. As we can see, when $\lambda_{\text{pres}}$ and $\lambda_{\text{attn}}$ vary within a reasonable range, our MagicTailor can consistently attain SOTA performance, revealing the robustness of MagicTailor on these crucial hyperparameters.
Table 5: Ablation of DS-Bal. We compare DS-Bal with potential variants, showing its excellence.
U-Net Variants CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO $\uparrow$ DreamSim $\downarrow$
Fixed ($\beta=0$) 0.268 0.850 0.803 0.293
Fixed ($\beta=1$) 0.270 0.851 0.808 0.286
Momentum ($\beta=0.5$) 0.268 0.850 0.805 0.290
Momentum ($\beta=0.9$) 0.269 0.850 0.808 0.288
Momentum (Ours) 0.270 0.854 0.813 0.279
4.5 Applications
Decoupled Generation.
After learning from a concept-component pair, MagicTailor can also enable decoupled generation. As shown in Fig. 7 (a), MagicTailor can generate the target concept and component separately in various and even cross-domain contexts. This should be credited to its remarkable ability to capture different-level visual semantics. Such an ability extends the flexibility of the possible combination between the concept and component.
Controlling Multiple Components.
In this paper, we focus on personalizing one concept and one component, because such a setting is enough to cover extensive scenarios, and can be further extended to reconfigure multiple components with an iterative procedure. However, as shown in Fig. 7 (b), our MagicTailor also exhibits the potential to control two components simultaneously. Handling more components remains a prospective direction of exploring better control over diverse elements for a single concept.
Figure 8: Ablation of loss weights. We report CLIP-T [11] for text alignment, and DreamSim [10] for identity fidelity as it is most similar to human judgments. For reference, we also present the second-best results in Table 1, highlighting our robustness.
Enhancing Other Generative Tools.
In Fig. 7 (c), we provide some interesting method combinations to show that our MagicTailor can enhance other generative tools. The combined tools include ContorlNet [56], CSGO [51], and InstantMesh [52]. As we can see, MagicTailor can be seamlessly integrated into these tools, furnishing them with an additional ability to control the concept’s component in their pipelines. For instance, working with MagicTailor, InstantMesh can conveniently achieve fine-grained design of 3D mesh, demonstrating the practicability of MagicTailor in collaborative applications.
5 Conclusion
We introduce component-controllable personalization, enabling precise customization of individual components within a concept. Our framework, MagicTailor, uses Dynamic Masked Degradation (DM-Deg) to suppress unwanted semantics and Dual-Stream Balancing (DS-Bal) to ensure balanced learning. Experiments show that MagicTailor sets a new standard in this task, with promising creative applications. In the future, we would like to extend our approach to broader image and video generation, enabling finer control over multi-level visual semantics for enhanced generative capabilities.
Supplementary Material
Appendix A More Details of Experimental Setup
A.1 Dataset
As there is no existing dataset specifically for component-controllable personalization, we curate a dataset from the internet to conduct experiments. Particularly, unlike previous works [36, 19] that focus on very few categories of concepts, the dataset contains concepts and components from various domains, such as characters, animation, buildings, objects, and animals. Overall, the dataset consists of 23 concept-component pairs totally with 138 reference images, where each concept/component contains 3 reference images and a corresponding category label. The scale of this dataset is aligned with the scale of those datasets used in the compared methods [11, 36, 19, 2, 38].
A.2 Implementation
We utilize Stable Diffusion 2.1 [35] as the pretrained T2I diffusion model. As commonly done, the resolution of reference images is set to 512 $\times$ 512. Besides, the rank and alpha of the LoRA module are set to 32. To simplify concept learning, we exclude the region of the target component from the segmentation masks of the target concept, e.g., remove the hair from the person in a “<person> + <hair>” pair. For the warm-up and DS-Bal stage, we set the learning rate to 1e-4 and 1e-5 and the training steps to 200 and 300. Moreover, the learning rate is further scaled by the batch size, which is set to completely contain a concept-component pair. For the cross-attention loss, we follow [2] to average the corresponding cross-attention maps at resolution 16 $\times$ 16 and normalized them to [0, 1]. The model is trained with an AdamW [23] optimizer and a DDPM [14] sampler. As done in [2], the tensor precision is set to float16 to accelerate training. For a fair comparison, all random seeds are fixed at 0 in each experiment, and all compared methods use the same implementation above except for method-specific configurations.
A.3 Evaluation
To generate images for evaluation, we carefully design 20 text prompts covering extensive situations, which are listed in Tab. 6. These text prompts can be divided into four aspects, including recontextualization, restylization, interaction, and property modification, where each aspect is composed of 5 text prompts. In recontextualization, we change the contexts to different locations and periods. In restylization, we transfer concepts into various artistic styles. In interaction, we explore the spatial interaction with other concepts. In property modification, we modify the properties of concepts in rendering, views, and materials. Such a group of diverse text prompts allows us to systemically evaluate the generalization capability of a method. We generate 32 images per text prompt for each pair, using a DDIM [41] sampler with 50 steps and a classifier-free guidance scale of 7.5. To ensure fairness, we fix the random seed within the range of [0, 31] across all methods. This process results in a total of 14,720 images for each method to be evaluated, ensuring a robust and thorough comparison.
Table 6: Text prompts used to generate evaluation images. These text prompts can be divided into four aspects: recontextualization, restylization, interaction, and property modification, covering extensive situations to systemically evaluate the generalization capability of a method. Note that “<placeholder>” will be replaced by the combination of pseudo-words (e.g., “<tower> with <roof>”) when generating evaluation images, and will be replaced by the combination of category labels (e.g., “tower with roof”) when calculating the metric of text alignment.
Recontextualization Restylization
“<placeholder>, on the beach” “<placeholder>, watercolor painting”
“<placeholder>, in the jungle” “<placeholder>, Ukiyo-e painting”
“<placeholder>, in the snow” “<placeholder>, in Pixel Art style”
“<placeholder>, at night” “<placeholder>, in Von Gogh style”
“<placeholder>, in autumn” “<placeholder>, in a comic book”
Interaction Property Modification
“<placeholder>, with clouds in the background” “<placeholder>, from 3D rendering”
“<placeholder>, with flowers in the background” “<placeholder>, in a far view”
“<placeholder>, near the Eiffel Tower” “<placeholder>, in a close view”
“<placeholder>, on top of water” “<placeholder>, made of clay”
“<placeholder>, in front of the Mount Fuji” “<placeholder>, made of plastic”
A.4 Automatic Metrics
We utilize four automatic metrics in the aspects of text alignment (CLIP-T [11]) and identity fidelity (CLIP-I [29], DINO [26], DreamSim [10]). To precisely measure identity fidelity, we improve the traditional measurement approach for personalization. This is because a reference image of the target concept/component could contain an undesired component/concept that is not expected to appear in evaluation images. Specifically, we use Grounded-SAM [34] to segment out the concept and component in each reference and evaluation image. Then, we further eliminate the target component from the segmented concept as we have done during training. Such a process is similar to the one adopted in [2]. As a result, using the segmented version of evaluation images and reference images, we can accurately calculate the metrics of identity fidelity.
A.5 User Study
We further evaluate the methods with a user study. Specifically, we design a questionnaire to display 20 groups of evaluation images generated by our method and other methods. Besides, each group also contains the corresponding text prompt and the reference images of the concept and component, where we adopt the same text prompts that are used to calculate CLIP-T. The results of our method and all the compared methods are presented on the same page. Clear rules are established for users to evaluate in three aspects, including text alignment, identity fidelity, and generation quality. Users are requested to select the best result in each group by answering the corresponding questions of these three aspects. We hide all the method names and randomize the order of methods to ensure fairness. Finally, 3,180 valid answers are collected for a sufficient quantitative evaluation of human preferences.
A.6 Compared Methods
In our experiments, we compare MagicTailor with SOTA methods in the domain of personalization, including Textual Inversion (TI) [11], DreamBooth-LoRA (DB) [36], Custom Diffusion (CD) [19], Break-A-Scene (BAS) [2], and CLiC [38]. We select these methods because TI, DB, and CD are three representatives of personalization frameworks and BAS and CLiC are highly relevant to learning fine-grained elements from reference images. For TI, DB, and CD, we use the third-party implementation in Diffusers [8]. For BAS, we use the official implementation. For CLiC, we reproduce it following the resource paper as the official code is not released. Unless otherwise specified, method-specific configurations are set up by following their resource papers or Diffusers. We empirically adjust the learning rate of CD and CLiC to 1e-4 and 5e-5 respectively, because they perform very poorly with the original learning rates. For a fair and meaningful comparison, these methods should be adapted to our task setting with minimal modification. Therefore, for those methods adopting a vanilla diffusion loss, we integrate the masked diffusion loss into them while using the same segmentation masks from MagicTailor.
Figure 9: Comparing with detailed-text-guided generation. We use GPT-4o to generate and merge detailed textual descriptions for the target concept and component, which are fed into Stable Diffusion 2.1 to conduct text-to-image generation. This paradigm cannot perform well and produce inconsistent images, while MagicTailor can achieve faithful and consistent generation.
Figure 10: Compatibility with different backbones. We equip MagicTailor with SD 1.5 [35], SD 2.1 [35], and SDXL [27]. The results show that MagicTailor can be generalized to multiple backbones, and a better backbone could provide better performance.
Appendix B Additional Comparisons
B.1 Detailed Text-Guided Generation
One might wonder if component-controllable personalization can be accomplished by providing detailed textual descriptions to the T2I model. To investigate this, we separately feed the reference images of the concept and component into GPT-4o [17] to obtain detailed textual descriptions for them. The text prompt we used is “Please detailedly describe the <concept/component> of the upload images in a parapraph”, where “<concept/component>” is replaced with the category label of the concept or component. Then, we ask GPT-4o to merge these textual descriptions using natural language, and input them into the Stable Diffusion 2.1 [35] to generate the corresponding images. Some examples for a qualitative comparison are shown in Fig. 9. As we can see, such an approach cannot achieve satisfactory results, because it is hard to guarantee that visual semantics can be completely expressed by using the combination of text tokens. In contrast, our MagicTailor is able to accurately learn the desired visual semantics of the concept and component from reference images, and thus lead to consistent and excellent generation in this tough task.
B.2 Commercial Models
It is also worth exploring whether existing commercial models, which can understand and somehow generate both text and images by themselves or other integrated tools, are capable of handling component-controllable personalization. We choose two widely recognized commercial models, GPT-4o [17] and Gemini 1.5 Flash [43], for a qualitative comparison. First, we separately feed the reference images of the concept and component into them, along with the text prompt of “The uploaded images contain a special instance of the <concept/component>, please mark it as #<concept/component>”, where “<concept/component>” is replaced with the category label of the concept or component. Then, we instruct them to perform image generation, using the text prompt of “Please generate images containing #<concept> with #<component>”, where “<concept>” and “<component>” are replaced with the category label of the concept and component, respectively. As shown in Fig. 12, these models struggle to reproduce the given concept, let alone reconfigure the concept’s component. Whereas, our MagicTailor achieves superior results in component-controllable personalization, using a dedicated framework designed for this task. It demonstrates that, even though large commercial models are able to tackle multiple general tasks, there is also plenty of room for the community to explore specialized tasks for real-world applications.
Figure 11: Effectiveness on difficult training pairs. We show the results of two hard cases involving large geometric discrepancy and cross-domain interactions, showing that MagicTailor can effectively handle such challenging scenarios.
Appendix C Generalizability Analysis
As discussed in Sec. A, we follow a strict and standardized setup to conduct performance evaluation for all the methods in our experiments. In this section, we perform a comprehensive generalizability analysis for MagicTailor by extending the setup to more flexible scenarios. Specifically, the analysis focuses on the model perspective, where we explore MagicTailor’s compatibility with different backbones (Sec. C.1), and the data perspective, where we study its effectiveness on difficult training pairs (Sec. C.2) and complex text prompts (Sec. C.3).
Figure 12: Comparing with commercial models. We input the reference images of the target concept and component to GPT-4o and Gemini, along with structured text prompts, for conducting image generation. Even though capable of handling multiple general tasks, these models still fall short in this task. In contrast, our MagicTailor performs well using a dedicated framework.
Figure 13: Effectiveness on complex text prompts. We present qualitative results generated with complex text prompts. In addition to those well-categorized text prompts, our MagicTailor can also follow more complex ones to generate text-aligned images.
C.1 Compatibility with Different Backbones
We use Stable Diffusion 2.1 (SD 2.1) [35], which is widely adopted by the compared methods, as our backbone in the main paper. It is worth noting that our MagicTailor is a general method also suitable for other T2I diffusion models, because only input images and the denoising paradigm are reformulated. To evaluate the compatibility, we employ MagicTailor in other backbones, including Stable Diffusion 1.5 (SD 1.5) [35] and Stable Diffusion XL (SDXL) [27]. As shown in Fig. 10, our MagicTailor can also achieve satisfactory results when cooperating with various backbones. Notably, we only change the LoRA rank for adapting to different learning capabilities, and directly use the original hyperparameter values, which shows the remarkable generalizability of MagicTailor.
C.2 Effectiveness on Difficult Training Pairs
Appropriate processing for reference images, such as filtering, scaling, and cropping, can help to construct a better training pair for personalization. Unfortunately, not all training pairs are ideal in real-world scenarios. Therefore, here we investigate the performance of MagicTailor when handling difficult training pairs. Specifically, we consider two hard cases. The first is a pair with large geometric discrepancy, such as “<person>” in an upper body portrait and “<hair>” in a profile photo. The second is a pair with cross-domain interactions, such as “<person>” and “<ear>” of dogs. As shown in Fig. 11, even facing these hard cases, MagicTailor is also capable of personalizing the target concept and component with high fidelity.
C.3 Effectiveness on Complex Text Prompts
In our experiments, we use well-categorized text prompts in four aspects, including recontextualization, restylization, interaction, and property modification, for qualitative and quantitative comparisons. Here we further evaluate MagicTailor’s performance on other complex text prompts involving more complicated contexts and elements. As shown in Fig. 13, our MagicTailor can also effectively generate text-aligned images when performing fidelity personalization, showing its ability to handle diverse user needs.
Figure 14: Ablation of DM-Deg via replacement with SID. We compare our DM-Deg with SID [18] that aims to produce informative prompts for training. Besides, we also present baseline (i.e., removing DM-Deg from MagicTailor) results for reference. This comparison indicates the effectiveness of our DM-Deg in addressing semantic pollution.
Appendix D Additional Ablation Studies
D.1 Effectiveness of DM-Deg over Using Informative Training Prompts
One might be curious about whether it is not necessary to employ the proposed DM-Deg, but perhaps to use informative text prompts during training to provide textual prior knowledge for learning the target concept and component. To investigate this, we use Selectively Informative Description (SID) [18] with GPT-4o [17] to construct text prompts for the target concept and component, and then use them for training. As shown in Fig. 14, such an approach cannot address semantic pollution well, where unwanted visual semantics still affect the personalized concept. In contrast, DM-Deg effectively prevents semantic pollution by dynamically perturbing those undesired visual semantics, verifying its remarkable significance in this task.
D.2 Necessity of Warm-Up Training
In MagicTailor, we start with a warm-up phase for the T2I model to preliminarily inject the knowledge for the subsequent phase of DS-Bal. Here we investigate the necessity of such a warm-up phase for generation performance. In Tab. 7, when removing the warm-up phase, even though MagicTailor could obtain slight improvement in text alignment, it severely suffers from the huge drop in identity fidelity. This is because such a scheme makes it difficult to construct a decent momentum denoising U-Net for DS-Bal. Whereas integrated with a warm-up phase, MagicTailor can achieve superior overall performance, which is attributed to the knowledge reserved from warm-up.
Table 7: Ablations of warm-up. We compare MagicTailor with the variant that removes warm-up. The results exhibit the significance of the warm-up stage for the framework of MagicTailor.
Warm-Up Variants CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO $\uparrow$ DreamSim $\downarrow$
w/o Warm-Up 0.272 0.844 0.793 0.320
w/ Warm-Up (Ours) 0.270 0.854 0.813 0.279
Figure 15: Ablation of the number of reference images. We present qualitative results to show that MagicTailor can still achieve satisfactory performance when provided only 1 or 2 reference image(s) per concept and component.
D.3 Performance on Different Numbers of Reference Images
As described in Appendix A, each concept/component contains 3 reference images in a concept-component pair of the dataset. Here we change the number of reference images to analyze the performance variation of MagicTailor. The qualitative results are presented in Fig. 15. When the number of reference images is reduced, MagicTailor can still show satisfactory performance. This also demonstrates that, while more reference images could lead to better generalization ability, one reference image per concept/component is enough to obtain a decent result with our MagicTailor.
D.4 Robustness on Linking Words
Generally, we use “with” to link the pseudo-words of the concept and component in a text prompt, e.g., “<person> with <beard>, in Von Gogh style”. Here we evaluate the robustness of our method on different linking words. We choose several words, which are commonly used to indicate ownership or association, to construct text prompts and then feed them into the same fine-tuned T2I model. As shown in Fig. 16, the generation performance of our MagicTailor remains robust regardless of the linking word used, exhibiting its flexibility to textual descriptions.
Figure 16: Ablation of linking words. We present qualitative results generated with different linking words, demonstrating the robustness of MagicTailor.
Appendix E More Qualitative Results
In Fig. 17 & Fig. 18, we provide more evaluation images for a substantial qualitative comparison. It can be clearly observed that semantic pollution remains an intractable problem for these compared methods. This is due to the leak of an effective mechanism to alleviate the T2I model’s perception for these semantics. To address this, our MagicTailor utilizes DM-Deg to dynamically perturb undesired visual semantics during the learning phase, and thus achieve better performance. On the other hand, the compared methods are also severely influenced by semantic imbalance, resulting in overemphasis or even overfitting on the concept or component. This is because the inherent imbalance of visual semantics complicates the learning process. In response to this issue, our MagicTailor applies DS-Bal to balance the learning of visual semantics, effectively showcasing its prowess in this tough task.
Figure 17: More qualitative comparisons. We present images generated by our MagicTailor and SOTA methods of personalization for various domains including characters, animation, buildings, objects, and animals. MagicTailor generally achieves promising text alignment, strong identity fidelity, and high generation quality.
Figure 18: More qualitative comparisons. We present images generated by our MagicTailor and SOTA methods of personalization for various domains including characters, animation, buildings, objects, and animals. MagicTailor generally achieves promising text alignment, strong identity fidelity, and high generation quality.